Use the given function f to find f(0) and solve f(x) = 0. Show work

f(x) = x^2 - x - 12
1
Expert's answer
2019-03-28T10:56:23-0400

1) Substitute 0 for x in the function:


"f(0)={{0}^{2}}-0-12=-12"

so


"f\left( 0 \right)=-12"

2) Solve

"f\left( x \right)=0"

We have equation

"{{x}^{2}}-x-12=0"

In order to solve this equation we use the quadratic formula. If 

"a{{x}^{2}}+bx+c=0"

then

"x=\frac{-b\pm \sqrt{{{b}^{2}}-4ac}}{2a}"

In our case

"a=1, b=-1, c=-12"

Substitute these numbers in the formula


"x=\frac{-\left( -1 \right)\pm \sqrt{{{\left( -1 \right)}^{2}}-4\cdot 1\cdot \left( -12 \right)}}{2\cdot 1}"

"x=\frac{1\pm \sqrt{1+48}}{2}"

"x=\frac{1\pm \sqrt{49}}{2}=\frac{1\pm 7}{2}"

"{{x}_{1}}=\frac{1+7}{2}=4,\,\,\,\,{{x}_{2}}=\frac{1-7}{2}=-3"

So our solution is


"{{x}_{1}}=4,\,\,{{x}_{2}}=-3"